OpenStudy (anonymous):

Evaluate and simplify tan 15º.

OpenStudy (anonymous):

You can write tan(15) as : \[\tan(45 - 30)\] Now use the formula: \[\tan(A - B) = \frac{\tan(A) - \tan(B)}{1 + \tan(A)\tan(B)}\]

OpenStudy (anonymous):

And : \[\tan(45) = 1\] \[\tan(30) = \frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

-sqrt3/2 is this correct

OpenStudy (anonymous):

No, check it once again..

OpenStudy (anonymous):

sqrt2/2 ?

OpenStudy (anonymous):

Are you guessing it??